OpenStudy (anonymous):

Sarah is 6 years older than 3 times the age of her son Tom. The difference in sarah age and Tom's age is 28 years. A. Write an equation using the information given? B. How old are Sarah and Tom

OpenStudy (anonymous):

anyone help?

OpenStudy (anonymous):

Its Algebra I guess and i need to do it for hw can anybody help please c:

OpenStudy (anonymous):

hmm anyone have a clue?

OpenStudy (anonymous):

ill help just give me a minute to work it out

OpenStudy (anonymous):

okay c:

OpenStudy (anonymous):

gimme 5 minutes im asking my teacher

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

There are two statements about the ages. That means there are TWO equations. Let S be Sarah's age and T be Tom's age. Then: S = 6 + 3T .... "Sarah is 6 years older than 3 times the age of her son Tom" |S - T| = 28 .... "The difference in Sarahs age and toms age is 28 years" Since you know from common sense that Sarah is older than Tom, you can drop the absolute value bars and have two equations in two unknows. The second is now: S - T = 28 .... which you can rewrite as: S = T + 28 .... and use this to subsitute for S in the first equation: (T + 28) = 6 + 3T T + 22 = 3T .... subtracted 6 from both sides 22 = 2T .... subtracted T from both sides T = 11 .... divide both sides by 2, and put the variable on the left Tom is 11 years old. Sarah's age is 6+3T = 6 + 3(11) = 6 + 33 = 39 years.

OpenStudy (anonymous):

so sarah is 39 years old?

OpenStudy (anonymous):

yes and tom is 11

OpenStudy (anonymous):

thanks so much c;

OpenStudy (anonymous):

ur welcome :)

OpenStudy (anonymous):

just ask if u have any problems in math

OpenStudy (anonymous):

i need to do a lot of these tyoe of problems for math and its super hard

OpenStudy (anonymous):

its pretty easy once u find the equation and yes i have some problems in math but i manage to get threw them haha what grade r u in?

OpenStudy (anonymous):

like 9 th

OpenStudy (anonymous):

i am too haha its hard huh?

OpenStudy (anonymous):

soorta

OpenStudy (anonymous):

u will get it soon enough haha i homeschool and i manage to get through it

OpenStudy (anonymous):

S = T(3) + 6 S - T = 28 the age difference 28 x's 3 + 6 = S S stands for Sara T stands for Tom

OpenStudy (anonymous):

i kinda already said that in the last message :D

OpenStudy (anonymous):

oh my bad, i didnt see that

OpenStudy (anonymous):

lol its totally fine :)